Let f(x) = |2x − 10| + 2 .
a) Show that f is continuous at x = 5.
b) Show that f is not differentiable at x = 5.
1
Expert's answer
2020-12-13T19:27:29-0500

a) A function "f(x)" is said to be continuous at a point "x=a" of its domain,iff "\displaystyle {\lim_{x\to a}}f(x)=f(a)"

"\implies \displaystyle {\lim_{x\to a^-}}f(x)=\displaystyle {\lim_{x\to a^+}}f(x)=f(a)"

We know that, "\displaystyle {\lim_{x\to a^-}}f(x)=\displaystyle {\lim_{h\to 0}}f(a-h)"

and "\displaystyle {\lim_{x\to a^+}}f(x)=\displaystyle {\lim_{h\to 0}}f(a+h)"

Here "f(x)=|2x-10|+2"

It can be written as,"f(x)=(2x-10)+2; when" "x>5"

"=-(2x-10)+2; when" "x<5"

"=2;when" "x=5"

According to the problem,

"\displaystyle {\lim_{x\to 5^-}}f(x)=\displaystyle {\lim_{h\to 0}}f(5-h)"

"=\displaystyle {\lim_{h\to 0}}[-[2(5-h)-10]]+2"

"=\displaystyle {\lim_{h\to 0}}(2h+2)"

"=2"

"\displaystyle {\lim_{x \to 5^+}}f(x)=\displaystyle {\lim_{h \to 0}}f(5+h)"

"=\displaystyle {\lim_{h\to 0}} [2(5+h)-10]+2"

"=\displaystyle {\lim_{h\to 0}}(2h+2)"

"=2"

And "f(5)=2"

"\therefore f(x)" is continuous at "x=5" .


b) A function f(x) is said to be differentiable at "x=a," iff "\displaystyle {\lim_{x\to a}}[f(x)-f(a)]/(x-a)" exists finite.

"\implies" "\displaystyle {\lim_{x\to a^-}}[f(x)-f(a)]/(x-a)=\displaystyle {\lim_{x\to a^+}}[f(x)-f(a)]/(x-a)"

"\implies \displaystyle {\lim_{h\to 0}}[f(a-h)-f(a)]/(-h)= \displaystyle {\lim_{h\to 0}}[f(a+h)-f(a)]/(h)"

Now,

"\displaystyle {\lim_{x\to 5^-}}[f(x)-f(5)]/(x-5)"

"=\displaystyle {\lim_{h\to 0}}[f(5-h)-f(5)]/(-h)"

"=\displaystyle {\lim_{h\to 0}}[(2h+2)-2]/(-h)"

"=\displaystyle {\lim_{h\to 0}}(-2)=-2"

"\displaystyle {\lim_{x\to 5^+}}[f(x)-f(5)]/(x-5)"

"=\displaystyle {\lim_{h\to 0}}[f(5+h)-f(5)]/(h)"

"=\displaystyle {\lim_{h\to 0}}[(2h+2)-2]/(h)"

"=\displaystyle {\lim_{h\to 0}}(2)=2"

"\therefore" "\displaystyle {\lim_{x\to 5^-}}[f(x)-f(5)]/(x-5)" "\neq" "\displaystyle {\lim_{x\to 5^+}}[f(x)-f(5)]/(x-5)"

Therefore "f(x)" is not differentiable at "x=5".